Which type of electromagnetic wave travels through
Answer #1

14 : option (A) is correct

15 option (A) is correct answer

16 from the Malus law

I1 = Io/2

I2 = I1 cos ^ theta

= Io/2 cos^2 (34 )

I2/Io = 34.4

(17)

convex mirror , meaning of that sentence is  the objects are made smaller by the mirror, so they appear farther away

As per guide lines I worked first 4 parts of the problem